please help me simplify the following:
a^-1-3a^-2/2^-2

Answer by benazir.sj@gmail.com(109)

a^-1-3a^-2/2^-2
solution:
(1/a -3/a^2)
____________
1/2^2

=(a^2 -3a)/a^3
_______________
1/4

=a(a-3)*4/a^3
=4(a-3)/a^2